# Multi-Objective min-max Online Convex Optimization

### v1

@@ 1.6 Our Contributions
    Compared to the discussed prior work, which has primarily focused on bounding the ‘respective’ regret under worst-case or adversarial inputs, establishing bounds with the adversarial input on the min–max static regret (5) appears to be substantially more challenging. We highlight the difficulty for one instance in Section 3, where even when full functions $f_{t},g_{t}$ are revealed before choosing action $x_{t}$, the greedy algorithm still has a linear regret.
@@
    Thus, we consider the min-max static regret problem in the i.i.d. setting where both $f_{t},g_{t}$ are chosen independently across $t$ with identical (unknown) distribution $\mathcal{D}$. Note that all the discussed applications and use cases for multi-objective OCO (e.g. fair machine learning, fair resource allocation, multi-task learning, network scheduling, multi-objective online ride-matching problem etc.) are well motivated in the i.i.d. case as well. Moreover, analytically also, the problem remains challenging with the i.i.d. input. From a theoretical advancement point of view similar approach was followed for the COCO problem where the earliest results were derived for the i.i.d. input Mahdavi et al. (2013), which were then later generalized to the worst case input.
    For the i.i.d. model, the expected cost of $\mathcal{A}$ is ${\bar{C}}_{\mathcal{A}}$ that is just the expected value of $C_{\mathcal{A}}$ in (2), while the $C_{\mathsf{OPT}}$ simplifies to
    ${\bar{C}}_{\mathsf{OPT}}=T\cdot\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x)\quad\text{with}\quad(x^{\star},\theta^{\star})=\arg\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x),$ (11)
    where $\mu(x)={\mathbb{E}}\{f_{t}(x),g_{t}(x)\}$, $\theta=(\theta_{1},\theta_{2})\in\Delta_{2}$ and $\Delta_{2}$ is the 2D2-D simplex. The metric is now the expected min-max static regret
    ${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}\{{\bar{C}}_{\mathcal{A}}-{\bar{C}}_{\mathsf{OPT}}\}.$ (12)
    Algorithm: For the i.i.d. model, the basic idea behind our solution approach for minimizing the expected min-max static regret (12) is as follows. From (11), the real goal for an algorithm is to find the optimal pair $(x^{\star},\theta^{\star})$. Since an online algorithm $\mathcal{A}$ does not know $\mathcal{D}$ ahead of time, finding $(x^{\star},\theta^{\star})$ is not possible. So we propose a simple algorithm that produces pair $(x_{t},\theta_{t})$ at time $t$ in an online manner that uses Hedge algorithm [^2] We describe the basic Experts problem and Hedge algorithm in Appendix E together with its regret guarantee for completeness purposes. Cesa-Bianchi and Lugosi (2006) to update $\theta_{t}$ to solve the inner maximization problem in (11) given $x_{t}$, and OGD to obtain actions $x_{t}$ given $\theta_{t}$ to solve the outer minimization (OCO) problem in (11) applied to the surrogate loss function $\theta_{t}^{\top}[f_{t}(x),g_{t}(x)]=\theta_{1t}f_{t}(x)+\theta_{2t}g_{t}(x)$ which is a proxy for $\theta^{\top}\mu(x)$, in an attempt to keep $(x_{t},\theta_{t})$ to stay close to $(x^{\star},\theta^{\star})$, so that the regret remains small.
    Note that (11) is not the usual online $\it min-max$} optimization problem Cardoso et al. (2019), where an online algorithm and adversary are playing against each other. Instead with (11) both $x$ and $\theta$ are algorithm’s choices depending on $\mathcal{D}$, that is under an adversarial control as defined in (12). The proposed alternating min-max algorithm is intuitive, however, the regret analysis is actually quite delicate, and heavily depends on our non-trivial decomposition of regret in (19) into three terms, which also motivates our algorithm.
    It is worth noting that there are several other candidate algorithms (e.g. primal-dual algorithm for COCO Yu et al. (2017) and OGD with surrogate cost function for COCO Sinha and Vaze (2024)) that appear to solve (11), however, analyzing them is rather challenging because of the outside $\max$ in the cost of an online algorithm in (5).
@@
    Guarantee: Using simple and elegant analysis via an insightful regret decomposition, we show that the expected min-max static regret (12) of the proposed algorithm is $O(\sqrt{T})$. This is similar to the short and elegant analysis of OGD algorithm’s Hazan (2019) performance for OCO in the adversarial input setting, that has been very useful for further progress. Even though we considered just two sequences $f_{t},g_{t}$ for simplicity, the result generalizes directly if there are $K$ sequences $f^{k}_{t},k=1,\dots,K$ and the expected min-max static regret (5) is $O(\sqrt{T\log K})$. See Remark 7 for more details.
@@
    Improved guarantee for GlobalExpertsProblem Even-Dar et al. (2009) in the i.i.d. input: As a special case of our result, we derive an improved regret guarantee of $O(\sqrt{T\log K})$ for the GlobalExpertsProblem Even-Dar et al. (2009) with the i.i.d. input compared to $O(\sqrt{T}\log K)$ of Even-Dar et al. (2009) which was applicable in the worst case/adversarial input setting.
@@
    With Bandit Feedback: We also extend our results to the bandit setting, where only function evaluations at the most recent actions are available without any gradient feedback. We show that the expected min-max static regret (12) is $O(T^{3/4}\sqrt{\log K})$ and $O(\sqrt{T\log K})$ for one-point and two-point bandit feedback, respectively, matching the best known bounds on OCO in the respective bandit settings.
1.7 Open Questions
  • We have only considered the i.i.d. input setting, and the major open question that remains is whether $o(T)$ min-max static regret (5) is achievable in the worst case input or not.
  • Lower bound on the expected min-max regret: Theoretically, deriving a lower bound on the expected min-max static regret matching the derived guarantee of $O(\sqrt{T})$ appears challenging, however, simulations (see Fig. 1) with $f_{t},g_{t}$ as random linear functions suggest that $\Theta(\sqrt{T})$ is the correct scaling for the expected min-max static regret.
    and the corresponding regret and CCV are
    respectively. It turns out that our proposed algorithm cannot be married readily with the prior best known approaches for COCO Sinha and Vaze (2024); Vaze and Sinha (2025); Sinha and Vaze (2025) for deriving simultaneous bounds on regret and CCV. Thus, how to solve the min-max OCO with constraints remains an interesting open problem.
@@ 2 Assumptions
We start by stating the assumptions that we will use throughout, that are standard in the OCO literature (Hazan, 2019).
@@
3 min-max OCO Problem (5) in the Adversarial Setting
In this section, we highlight the difficulty in solving the min-max OCO Problem (5) in the adversarial input setting, where the input can be chosen adversarially depending on the choice of an online algorithm.
To illustrate the basic difficulty, we even allow the algorithm to know $f_{t},g_{t}$ completely before choosing its action $x_{t}$ at time $t$. Under this enhanced information setting, consider the greedy (locally optimal) algorithm that chooses
$x_{t}=\min_{x\in\mathcal{X}}\max\{f_{t}(x),g_{t}(x)\}.$
Recall that when $g_{t}=f_{t}\ \forall\ t$, i.e. in the OCO setting, this algorithm will have non-positive regret $R^{\text{s}ingle}_{\mathcal{A}}(T)$ (1) since $f_{t}(x_{t})\leq f_{t}(x)$ for any $x$. However, as we show next, the greedy algorithm suffers a linear min-max static regret (5).
Let $T=2N$ and consider the following linear functions on $\mathcal{X}=[0,1]$, for $j=1,\dots,N$
@@
Clearly,
$\arg\min_{x\in\mathcal{X}}\{\max(f_{2j-1}(x),g_{2j-1}(x))\}=1$ $\quad\text{and}\quad\arg\min_{x\in\mathcal{X}}\{\max(f_{2j}(x),g_{2j}(x))\}=0.$
Thus, the greedy algorithm chooses $x_{2j-1}=1$ and $x_{2j}=0,$ and the cost $C_{\mathcal{A}}$ (2) of the greedy algorithm is
$\max\left\{\sum_{j=1}^{N}\left(f_{2j-1}(x_{2j-1})+f_{2j}(x_{2j})\right),\sum_{j=1}^{N}\left(g_{2j-1}(x_{2j-1})+g_{2j}(x_{2j})\right)\right\}$
$=\max\left\{\sum_{j=1}^{N}\left(f_{2j-1}(1)+f_{2j}(0)\right),\sum_{j=1}^{N}\left(g_{2j-1}(1)+g_{2j}(0)\right)\right\}=\max\left\{\sum_{j=1}^{N}(1+0),\sum_{j=1}^{N}(1+0.8)\right\}=1.8N.$
However, if an algorithm chooses a single static action $x^{\star}=0$ for all times, its cost $C_{\mathcal{A}}$ (5) is $1.2N$, and hence
$C_{\mathsf{OPT}}\leq 1.2N.$
Thus, the min-max static regret (5) of the greedy algorithm is linear in $T$. This example highlights the key difficulty in minimizing the min-max static regret in the adversarial/worst case input. We note that this example does not preclude more intelligent algorithms that take a global view in trying to balance the two costs, such as primal-dual, OGD+Hedge or FTRL to achieve sub-linear min-max static regret (5), however, the stringent min-max static regret definition makes the analysis highly challenging. We will make additional remarks on the technical difficulty in Appendix F. Thus, from here on we consider the i.i.d. input that is defined next.
4 i.i.d. Input model
We consider the i.i.d. input model, where at time $t$, loss functions $(f_{t},g_{t})$ are chosen independently and are identically jointly distributed as $\mathcal{D}$ (unknown) across all $t$. We assume that Assumption 1 and 2 holds, and consider the same information structure as defined in Assumption 3.
Let $\lambda_{t}=(f_{t},g_{t})$ and recall that $\Delta_{2}$ is the 2D2-D simplex, $\theta=(\theta_{1},\theta_{2})\in\Delta_{2}$, and
$\Lambda_{t}(x,\theta):=\theta_{1}f_{t}(x)+\theta_{2}g_{t}(x).$
Following (2), the expected cost of any online algorithm $\mathcal{A}$ with the i.i.d. input which chooses actions $x_{t}$ is
${\bar{C}}_{\mathcal{A}}={\mathbb{E}}\left\{\max\Big\{\sum_{t=1}^{T}f_{t}(x_{t}),\sum_{t=1}^{T}g_{t}(x_{t})\Big\}\right\}={\mathbb{E}}\left\{\max_{\theta\in\Delta_{2}}\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta)\right\}.$ (16)
while from (11)
${\bar{C}}_{\mathsf{OPT}}=T\cdot\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x)\quad\text{with}\quad(x^{\star},\theta^{\star})=\arg\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x),$ (17)
where $\mu(x)={\mathbb{E}}\{f_{t}(x),g_{t}(x)\}$.
Consequently, the expected min-max static regret ${\bar{R}}_{T}(\mathcal{A})$ following (12) is
${\bar{R}}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}\left\{{\bar{C}}_{\mathcal{A}}-{\bar{C}}_{\mathsf{OPT}}\right\}.$ (18)
We refer to (18) as the i.i.d. min-max OCO problem and for the rest of the paper, consider it with the goal of deriving an algorithm that has $o(T)$ expected min-max static regret.
Remark 3.
Note that even i.i.d. min-max OCO problem is theoretically challenging, and no prior results apply directly.
@@ 4.1 Algorithm
In this section, we present an algorithm to solve the i.i.d. min-max OCO problem (18). Towards that end, if suppose an online algorithm knew the optimizing $\theta^{\star}$ in (17), then by executing an OGD algorithm for OCO with a single function $\theta_{1}^{\star}f_{t}+\theta_{2}^{\star}g_{t}$ at time $t+1$, it would achieve a regret of $O(\sqrt{T})$ using standard results Hazan (2019). However, the optimizing $\theta^{\star}$ that depends on $\mathcal{D}$ (unknown) remains unknown, but, this observation motivates our algorithm that constructs
i) $\theta_{t}$ at time $t$ as a surrogate for $\theta^{\star}$ using the Hedge algorithm (gains version) treating the two function sequences $\{f_{t}\}_{t=1}^{T}$ and $\{g_{t}\}_{t=1}^{T}$ as the gains of the two experts (see Section E for basic definition of the Experts Problem and the Hedge algorithm), and
Algorithm 1 Minimization of the Maximum of The Sum of Two Convex Loss Sequences
@@
5:  Using current $\theta_{t}$, form convex function $x\mapsto{\tilde{\Lambda}}_{t}(x,\theta_{t}):=\theta_{t,1}f_{t-1}(x)+\theta_{t,2}g_{t-1}(x)$
6:  OGD: Gradient step: $y_{t}=x_{t-1}-\eta_{x,t}\nabla{\tilde{\Lambda}}_{t}(x_{t-1},\theta_{t})$ and find $x_{t}=\text{Proj}_{\mathcal{X}}(y_{t})$
7:  Play action $x_{t}$
8:  Evaluate $\lambda_{t}=(f_{t}(x_{t}),g_{t}(x_{t}))$
9:  Hedge: Update weights: $w_{t+1,i}=w_{t,i}\exp(\eta_{\theta,t}\lambda_{t,i})$ for $i=1,2$
10:  Normalize: $\theta_{t+1}=w_{t+1}/\|w_{t+1}\|_{1}$
11:end for

### v3

@@ 2.5 Our Contributions
    We begin by deriving a lower bound of $\Omega(T)$ on the min-max static regret (3) of any online algorithm with adversarial input. The input construction is inspired from Si Salem et al. (2023) and is modified to suit our specific problem.
@@
    Motivated by the impossibility of sublinear min-max static regret in the adversarial setting, we consider a stochastic i.i.d. model, [^2] We also consider the Martingale difference and Markov input models in Appendix N and O. where loss pairs $(f_{t},g_{t})$ are drawn from a fixed unknown distribution $\mathcal{D}$. This setting is standard for applications like fair resource allocation and machine learning and remains analytically challenging.
    Under the i.i.d. model, the benchmark is ${\mathbb{E}}\{C_{\mathsf{OPT}}\}$ where $C_{\mathsf{OPT}}$ has been defined in (2). Directly analyzing the min-max static regret (3) with respect to ${\mathbb{E}}\{C_{\mathsf{OPT}}\}$ is technically challenging, since $C_{\mathsf{OPT}}$ depends on the entire realized sequence of losses. Instead, we introduce a deterministic surrogate benchmark $\bar{C}_{\mathsf{OPT}}$ obtained by averaging out the randomness in the losses:
    $\bar{C}_{\mathsf{OPT}}=T\cdot\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x),\qquad\mu(x)={\mathbb{E}}\{[f_{t}(x),g_{t}(x)]\},$ (7)
    where $\Delta_{K}$ is the $K$-dimensional simplex. Importantly, as shown in Appendix D,
    ${\mathbb{E}}\!\left\{\,\big|C_{\mathsf{OPT}}-\bar{C}_{\mathsf{OPT}}\big|\,\right\}=O\!\left(\sqrt{T\log T}\right).$
    Using $\bar{C}_{\mathsf{OPT}}$ as the benchmark, we define the expected min-max static regret of an online algorithm $\mathcal{A}$ as
    $\bar{R}_{\mathcal{A}}(T)=\sup_{\mathcal{D}}\left\{{\mathbb{E}}\{C_{\mathcal{A}}\}-\bar{C}_{\mathsf{OPT}}\right\}.$ (8)
    For clarity, we refer to regret measured with respect to ${\mathbb{E}}\{C_{\mathsf{OPT}}\}$ as the true min-max static regret, and regret measured with respect to $\bar{C}_{\mathsf{OPT}}$ as the expected min-max static regret.
    In the remainder of the paper, we bound $\bar{R}_{\mathcal{A}}(T)$. From the above concentration bound, the true min-max static regret satisfies ${\mathbb{E}}\{C_{\mathcal{A}}\}-{\mathbb{E}}\{C_{\mathsf{OPT}}\}\;\leq\;\bar{R}_{\mathcal{A}}(T)+O\!\left(\sqrt{T\log T}\right)$.  [^4] With $K$ sequences, the additive term becomes $O\!\left(\sqrt{T\log(TK)}\right)$.
    Even under the i.i.d. model, the lower bound on the true min–max static regret is $\Omega(\sqrt{T})$. This follows by taking $f_{t}\equiv g_{t}$ for all $t$, which reduces the problem to standard OCO, for which ${\mathbb{E}}\{R^{\text{single}}_{\mathcal{A}}(T)\}=\Omega(\sqrt{T})$ even under i.i.d. inputs (e.g., random linear losses; see Hazan (2019)).
@@
    Algorithm While considering expected min-max static regret (8), the structure of (7) reveals the core intuition for our approach: the optimal policy corresponds to a saddle point $(x^{\star},\theta^{\star})$ that solves ${\bar{C}}_{\mathsf{OPT}}$ in (7). Since the distribution $\mathcal{D}$ is unknown, we cannot compute this pair offline. Instead, we propose a natural online algorithm (Algorithm 1) that interlaces two classic strategies: i) Hedge to update the weights $\theta_{t}$ (solving the inner maximization over objectives in (7)), and ii) OGD to update the action $x_{t}$ (solving the outer minimization on the surrogate loss $\theta_{t}^{\top}[f_{t},g_{t}]$ in (7)).
    While (7) resembles a zero-sum game Cardoso et al. (2019), the problem is fundamentally different: here, the learner controls both $x_{t}$ and $\theta_{t}$ to track an unknown distribution. The novelty of our approach lies not in the individual choice of Hedge or OGD, but in their principled combination which simplifies an otherwise intractable analysis. While other candidates like FTRL or algorithms for COCO Yu et al. (2017); Sinha and Vaze (2024) have an intuitive appeal for minimizing (8), there are obvious analytical roadblocks caused by the global $\max$ in the cost function (1) even with stochastic i.i.d. input. Our framework resolves this difficulty through a modular regret decomposition and decouples the dynamics of Hedge and OGD that renders the proof both modular and elegant.
    We prove that our algorithm achieves an expected min-max static regret (8) of $O(\sqrt{T})$ (or $O(\sqrt{T\log(K)})$ for $K$ objectives). Analogous results for Martingale difference and Markov input models are provided in Appendix N and O.
    Recall that the $\alpha$-fair model of Si Salem et al. (2023) is equivalent to our min-max setting when $\alpha\rightarrow\infty$. Thus, importantly, compared to the $O(\alpha\sqrt{T})$ regret bound of Si Salem et al. (2023) we obtain an $\alpha$-independent regret bound.
@@
    With Bandit Feedback: Exploiting our modular proof approach, we easily extend our results to the bandit setting, where only function evaluations at the most recent actions are available without any gradient feedback. We show that the expected min-max static regret (8) is $O(T^{3/4}\sqrt{\log K})$ and $O(\sqrt{T\log K})$ for one-point and two-point bandit feedback, respectively, matching the best known bounds on OCO in the respective bandit settings.
@@
    We also consider the min-max generalization of the universal prediction of binary sequences Feder et al. (2002); Cover (1966), where at each time $t$, an online algorithm $\mathcal{A}$ has to choose a binary action $x_{t}$ before two binary sequences $b_{t}^{1},b_{t}^{2}$’s are revealed. The cost of $\mathcal{A}$ on sequence $k$ is its Hamming distance between $\{x_{t}\}_{t=1}^{T}$ and $\{b_{t}^{k}\}_{t=1}^{T},k=1,2$, and the overall cost of $\mathcal{A}$ is the maximum of its cost on the two sequences. Even for this case, we show that the min-max static regret of any online algorithm is $\Omega(T)$ with adversarial input. With stochastic input where $(b_{t}^{1},b_{t}^{2})$ are i.i.d. across time, we propose an algorithm with $O(1)$ true min-max static regret. In comparison, for the classical version Feder et al. (2002); Cover (1966) with only a single sequence $b_{t}^{1}$, Follow-the-Leader (FTL) algorithm achieves $O(1)$ regret when the input is i.i.d. Feder et al. (2002), while Cover (1966) achieves $O(\sqrt{T})$ regret in the adversarial case.
We discuss the applications of multi-objective OCO and research questions left open in Appendix A, while numerical results are provided in Appendix 10.
@@ 3 Assumptions
We make the following assumptions that are standard in the OCO literature Hazan (2019).
@@
4 Warmup
In this section, we highlight the difficulty in solving the min-max OCO Problem (3) in the adversarial input setting. To illustrate the basic difficulty, we even allow an online algorithm to know $f_{t},g_{t}$ completely before choosing its action $x_{t}$ at time $t$. Under this enhanced information setting, consider the greedy (locally optimal) algorithm that chooses $x_{t}=\min_{x\in\mathcal{X}}\max\{f_{t}(x),g_{t}(x)\}.$ Recall that when $g_{t}=f_{t}\ \forall\ t$, i.e. in the OCO setting, this algorithm will have non-positive regret $R^{\text{s}ingle}_{\mathcal{A}}(T)$ since $f_{t}(x_{t})\leq f_{t}(x)$ for any $x$. However, as we show next, the greedy algorithm suffers a linear min-max static regret (3). Let $T=2N$ and for $j=1,\dots,N$, consider the following linear functions on $\mathcal{X}=[0,1]$,
@@ 4 Warmup
Clearly, $\arg\min_{x\in\mathcal{X}}\{\max(f_{2j-1}(x),g_{2j-1}(x))\}=1\ \text{and}\arg\min_{x\in\mathcal{X}}\{\max(f_{2j}(x),g_{2j}(x))\}=0.$
Thus, the greedy algorithm chooses $x_{2j-1}=1$ and $x_{2j}=0,$ and the cost $C_{\mathcal{A}}$ (1) of the greedy algorithm is $1.8N$. However, if an algorithm $\mathcal{A}_{1}$ chooses a single static action $x^{\star}=0$ for all times, its cost $C_{\mathcal{A}_{1}}$ (3) is $1.2N$, and hence $C_{\mathsf{OPT}}\leq 1.2N.$ Thus, the min-max static regret (3) of the greedy algorithm is linear in $T$.
5 Lower bound on min-max static regret (3) for all online algorithms
We formalize the intuition presented in Section 4 as follows.
Theorem 4.
The min-max static regret (3) for any online algorithm is $\Omega(T)$.
The proof of Theorem 4 can be found in Appendix C. The lower bound derived in Theorem 4 uses the fact that both $f_{t}$ and $g_{t}$ are convex and not strongly convex. With strongly convex functions, the lower bound may weaken. In light of Theorem 4, next, we consider the stochastic i.i.d. case for solving the min-max OCO problem which is also non-trivial.
6 Stochastic i.i.d. input model
We consider the i.i.d. input model, where at time $t$, loss function pair $(f_{t},g_{t})$ is chosen independently and is identically (jointly) distributed as $\mathcal{D}$ (unknown) across all $t$. The objective is to minimize the expected min-max static regret (8).
@@ 6.1 Algorithm
In this section, we present an algorithm to solve the i.i.d. min-max OCO problem (8).
Algorithm 1 Hedge+OGD
@@
5:  Form convex function $x\mapsto{\tilde{\Lambda}}_{t}(x,\theta_{t}):=\theta_{t,1}f_{t-1}(x)+\theta_{t,2}g_{t-1}(x)$
6:  OGD: $y_{t}=x_{t-1}-\eta_{x,t}\nabla{\tilde{\Lambda}}_{t}(x_{t-1},\theta_{t})$ and find $x_{t}=\text{Proj}_{\mathcal{X}}(y_{t})$
7:  Play action $x_{t}$  Evaluate $\lambda_{t}=(f_{t}(x_{t}),g_{t}(x_{t}))$
8:  Hedge: $w_{t+1,i}=w_{t,i}\exp(\eta_{\theta,t}\lambda_{t,i}),i=1,2$
9:  Normalize $\theta_{t+1}=w_{t+1}/\|w_{t+1}\|_{1}$
10:end for
Towards that end, if suppose an online algorithm knew the optimizing $\theta^{\star}$ in (7), then by executing an OGD algorithm for OCO with a single function $\theta_{1}^{\star}f_{t}+\theta_{2}^{\star}g_{t}$ at time $t+1$, it would achieve a regret of $O(\sqrt{T})$ using standard results Hazan (2019). However, the optimizing $\theta^{\star}$ that depends on $\mathcal{D}$ (unknown) remains unknown, but, this observation motivates our algorithm that constructs
i) $\theta_{t}$ at time $t$ as a surrogate for $\theta^{\star}$ using the Hedge algorithm (gains version) treating the two function sequences $\{f_{t}\}_{t=1}^{T}$ and $\{g_{t}\}_{t=1}^{T}$ as the gains of the two experts (see Section J for basic definition of the Experts Problem and the Hedge algorithm), and